OpenStudy (anonymous):

Jane's 85 on her english test is 37 points less than twice her grade on a science test. Write an equation to represent this.

OpenStudy (anonymous):

i put

OpenStudy (anonymous):

85=37-x2

OpenStudy (anonymous):

x being her science test score

OpenStudy (anonymous):

2x-37=85

OpenStudy (anonymous):

it isnt the same thing?

OpenStudy (anonymous):

No...85 is 37points less than 2x Your answer meant 85 is 2x points less than 37, which is wrong

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

2. three less than twice a number js 9

OpenStudy (anonymous):

2x-3=9?

OpenStudy (anonymous):

exactly..Well done!!!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

3. four less than the quotient of a number and 8

OpenStudy (anonymous):

4-n/8

OpenStudy (anonymous):

4 less than that quotient is asked..Try again

OpenStudy (anonymous):

the other way around

OpenStudy (anonymous):

x less y is y-x

OpenStudy (anonymous):

n/8-4?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

lol am i annoying you?

OpenStudy (anonymous):

Thats okay...

OpenStudy (anonymous):

5. choose the situation matching the equation 12m+25=350

OpenStudy (anonymous):

a.a game system costs $350 . andrew is putting a $12 deposit down and then paying $25 per month until it is payed it off

OpenStudy (anonymous):

b.april is training for a swim meet. she swam 25 laps on saturday every other day she will swim 12 laps until she reached her goal of 350

OpenStudy (anonymous):

Sandra has 12 months to plan her wedding. she made a $25 payment to the potographer who will charge her $350

OpenStudy (anonymous):

okay its B

OpenStudy (anonymous):

Joe collects baseball cards. he has a card worth$25 he adds 12 cards per week until he has a total of 350 cards

OpenStudy (anonymous):

Why is it b?

OpenStudy (anonymous):

april had to swim 350laps. She swam 25laps on saturday and 12laps each day to complete 350laps. So, 350 = 25 +12m (m is the number of days after saturday)

OpenStudy (anonymous):

..reading..

OpenStudy (anonymous):

ok. tnx:)

OpenStudy (anonymous):

i yhink i can do the rest. tnx:)

OpenStudy (anonymous):

oh that's good!!!

OpenStudy (anonymous):

lol:)

OpenStudy (anonymous):

tnx:)

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

@nincompoop It was asked just to represent the equation and not to find the variable!!! :D

OpenStudy (anonymous):

o i do have 1 more

OpenStudy (anonymous):

sure ask

OpenStudy (anonymous):

Olive wants to share her 48 cookies with her 3 friends.write an expression showing how many cookes each friend will get

OpenStudy (anonymous):

48/3 = 16

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

Each of her friends get 16 cookies

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

write a phrase for 3x-7

OpenStudy (anonymous):

7 less than thrice of x

OpenStudy (anonymous):

seven less thsn three times a numbee?

OpenStudy (anonymous):

number*

OpenStudy (anonymous):

that's also right...

OpenStudy (anonymous):

k tnx

OpenStudy (anonymous):

a number increased by ninety three

OpenStudy (anonymous):

x + 93

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

Okay i gotta go to bed bye GOOD NIGHT(for me though!)

OpenStudy (anonymous):

:) k bye